OpenStudy (anonymous):

I am struggling with this, Prove the theorem that if limx→d h(x) = P and limx→d k(x) = Q and h(x) ≥ k(x) for all x in an open interval containing d, then P ≥ Q by using the formal definition of the limit, showing all work.

OpenStudy (anonymous):

Suppose Q<P, let \[ a= \frac {P-Q}3 \] We can find an x close to d so that \[ |h(x)-P| < \frac a 3 \\ |k(x)-Q| < \frac a 3 \\ \] This imples that \[ k(x) < h(x) \] A contradiction which implies what we want.

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

YW